OpenStudy (waheguru):

you pay 60 dollars including tax and tax was 13percent what was the price b4 taxes

OpenStudy (cwrw238):

113 % = 60 dollars 100% = (60 * 100) / 113

OpenStudy (cwrw238):

this comes to $53.10

OpenStudy (waheguru):

please explain more

OpenStudy (waheguru):

why 60 x100?

OpenStudy (waheguru):

..........................

OpenStudy (cwrw238):

its a matter of proportion: 60 dollars equals the original price + 13% tax = 113% of original price and original price is 100% so 113% = 60 dollars 1% = 60 / 113 dollars 100% = (60/113) * 100